Three individuals form a partnership and agree to divide the profits equally x invests 9,000, y invests 7,000 z invests 4,000. If the profits are 4,800, how much less does x recieve compared to having the profits divided into portions to the amounts xyz
Answer by Mathtut(3670)

9000+7000+4000=20000
:
if profits were proportionate to the amounts invested then
:
x would receive 9/20 of the profit
:
9/20(4800)=2160
:
split evenly x will receive 4800/3=1600
:
the difference is 2160-1600=$560 dollars
:
so x is receiving 560 less than if the profits per proportionate to $$ invested